OpenStudy (ttp):

1kg is equal to 100 g 10000dg 1000 mg 1000 cg

OpenStudy (kc_kennylau):

firstly you have to know the prefixes by heart: k=10^3 d=10^-1 m=10^-3 c=10^-2

OpenStudy (kc_kennylau):

100g=100g 10000dg=10000*0.1g=1000g 1000mg=1000*0.001g=1g 1000cg=1000*0.01g=10g Therefore B

undeadknight26 (undeadknight26):

I agree with him...

OpenStudy (ttp):

Thanks!

OpenStudy (kc_kennylau):

no problem :)